# Tie-Breaking Rule Based on Partial Proof of Work in a Blockchain

### v1

@@ Abstract
Numerous methods have been proposed for suppressing intentional forks by attackers in blockchain systems. Among these, last-generated rules, which select the latest chain among chains in a tie, are effective methods that do not require significant changes to the blockchain protocol. However, existing methods either require a trusted third party or rely on timestamps that attackers can manipulate which makes applying a last-generated rule to existing systems such as Bitcoin challenging.
To address these issues, we propose a last-generated rule that can be easily applied to existing proof of work blockchain systems. Our method uses partial proof of work, which does not function as a block, as a time standard with finer granularity. Only weak synchronization, which is already met by existing systems, is required for effective functioning.
@@
Index Terms:
blockchain, fork choice rule, tie-breaking rule, selfish mining, partial proof of work
@@ I Introduction
In blockchain systems, transactions are processed in units known as blocks. Each block consists of a header and list of transactions that are processed in that block. A block is valid when its header’s hash value is less than a certain target value. The process of identifying such a header is known as mining. The nodes that participate in the system and generate blocks are referred to as miners. When miners successfully generate a block, they publish it and the published block is shared across the network. Miners can receive rewards for the blocks that they successfully generate.
@@
The security of a blockchain system depends on each miner adhering to the aforementioned protocol. Therefore, it is desirable that following the protocol is the most profitable action for each miner. However, Eyal et al. introduced a new mining strategy known as selfish mining (SM) and demonstrated that blockchain does not always satisfy the necessary incentive compatibility[3]. Specifically, they showed that miners can gain more rewards by withholding a successfully generated block instead of publishing it immediately, and then releasing it when another node publishes a block.
Several studies have been conducted on suppressing such unintended behaviors [3, 4, 5, 6, 7, 8, 9]. For instance, Eyal et al. proposed a tie-breaking rule that randomly selects a chain during chain ties (hereafter referred to as the random rule), thereby introducing the approach of modifying tie-breaking rules as a countermeasure to SM.
Additionally, to the best of our knowledge, Heilman was the first to propose a tie-breaking rule that selects the latest chain (hereafter referred to as last-generated rules)[4]. Last-generated rules enable each miner to distinguish between chains that are withheld by attackers and others. Consequently, last-generated rules can suppress intentional forks by attackers more effectively than the random rule.
@@
The proposed method is practical and applicable to existing systems that employ PoW blockchains, including Bitcoin. Our proposed method requires only that there exists a widely known upper bound on the message propagation times, and there exists a widely known upper bound on the absolute values of clock drifts. Our approach does not require a specific trusted third party. Furthermore, it does not require updates that can destabilize the system, such as soft forks or hard forks.
@@
The remainder of this paper is organized as follows. Section II describes intentional forks by attackers. Section III discusses related works. Section IV describes the proposed method. Section V demonstrates the effectiveness of the proposed method. Section VI presents a detailed discussion, primarily focusing on the application of the proposed method and consideration of clock drift. Section VII concludes the paper.
@@ II Intentional Forks by Attackers
In a blockchain, a fork refers to a situation in which a chain splits into multiple branches. Two types of forks exist: those caused intentionally by miners not following the protocol and those that occur naturally. Attacks such as SM[3, 8, 10, 11, 12] are typical examples of intentional fork exploitation. Among intentional forks, chain ties that are intentionally created by attackers are known as intentional chain ties. SM and similar attacks exploit these intentional chain ties. The proposed method is a tie-breaking rule that aims to suppress intentional chain ties. Suppressing intentional chain ties leads to suppressing intentional forks, which in turn helps to mitigate attacks such as SM.
We consider SM as an example to demonstrate the exploitation of intentional forks. First, the proportion of a miner’s own blocks in the main chain is defined as the relative revenue. The goal of SM is to achieve a higher relative revenue than the proportion of the attacker’s hashrate in the total network. If this is accomplished, the attack is considered successful. The attacker withholds the block that they have successfully mined instead of publishing it immediately. When an honest miner generates a block, the attacker publishes their block, creating a chain tie. If a block is generated that extends the attacker’s chain after the chain tie, the honest miner’s block is invalidated, and vice versa. If the attacker mines blocks consecutively, they publish their block when the length difference between their chain and the honest miner’s chain is one. In this scenario, the honest miner’s chain is invalidated because the attacker’s chain is longer. Thus, the attacker can invalidate the blocks that are created by honest miners and increase their own relative revenue by manipulating the timing of publishing their mined blocks in a manner that does not follow the protocol.
@@
Match against pre-generated block is difficult under the first-seen rule but becomes feasible under last-generated rules. Last-generated rules impose a time limit on accepting chains in a tie to address this issue. This time limit is referred to as the acceptance window $w$. Without such a limit, attackers can always initiate chain ties against honest miners and invalidate their blocks. Conversely, Match against pre-generated block can be mitigated to an extent by setting this limit. We detail the extent to which setting an acceptance window quantitatively suppresses Match against pre-generated block in Section V-B.
@@ III Related Work
All of the aforementioned methods are tie-breaking rules. Zhang et al. proposed a fork choice rule known as weighted FRP to suppress SM[6]. This method involves including uncle blocks in a block to consider the number of published blocks in the main chain selection. However, this approach requires changes to the fork choice rule, which may be expensive to implement and can result in multiple forks. In addition, this method may lead to prolonged forks in situations in which the network is fragmented. While this method is a fork choice rule, it does not specifically address which action should be taken in the case of chain ties. The proposed method is a tie-breaking rule and can also be applied to this method.
Solat et al. proposed a method named ZeroBlock in which empty blocks are generated when no block is produced for a certain period, thereby invalidating withheld blocks[7]. However, this method requires precise clock synchronization across the entire system, which is a stringent requirement, especially for global-scale decentralized systems such as Bitcoin. Moreover, this approach assumes that blocks in PoW blockchains are generated at regular intervals, but the interval between block generations follows an exponential distribution in actual systems, which would limit the effectiveness of their proposed method. Additionally, implementing this method requires significant updates such as hard forks, which makes it challenging to apply to systems that are already operational (e.g., Bitcoin).
Saad et al. proposed a fork choice rule with a similar concept to that of Lee et al.[8]. Their method uses high accuracy in predicting the block height, which is the time at which transactions are likely to be included in a block. Similar to the method of Lee et al., the average predicted processing times of transactions that are included in a block are used to distinguish between blocks generated by attackers and honest miners. However, their method also does not address the possibility of attackers including arbitrary transactions in their blocks. Furthermore, it is applicable only to specific patterns of attacker behavior, which limits its effectiveness.
@@ IV Proposed Method
Our proposed method, which is similar to that proposed by Heilman, is a type of last-generated rule. We use partial PoW, which originally had no functional use as a block, as a new time standard. Partial PoWs are easier to generate and can be produced in larger quantities than in blocks, and thus can function as a more granular clock.
Widespread sharing of partial PoWs across the system is required for our method to be effective. However, a partial PoW typically has a smaller size than blocks and does not consume significant network resources of miners. In addition, the implementation of our method has a negligible impact on the transaction-processing capacity of the system. Our method does not require updates that could divide the hashrate of the system, such as hard forks or soft forks, nor does it require a trusted third party. Although a certain level of synchrony in the message propagation times and the progression of miners’ clocks (clock drift) are necessary, these requirements are achievable in most blockchain systems and do not compromise the high level of asynchrony that is inherent in PoW blockchains. These factors make our method more practical than previous methods.
The proposed method is a tie-breaking rule for resolving chain ties and can be applied to existing fork choice rules. For example, it can be applied to rules such as GHOST[15], which had been partially adopted in Ethereum[13], and weighted FRP[6].
@@ IV-A Our Model
The final assumption is that the impact of clock drift among miners is sufficiently small to be negligible. For simplicity, We adopt this assumption in our explanation. However, this assumption can easily be replaced with a weaker one: there exists a widely known value $D$ such that the absolute value of the clock drift of any miner is less than $D$. We elaborate on how clock drift affects the proposed method in Section VI-D. In Bitcoin, the PoW blockchain ensures functionality provided that the clock discrepancies of miners do not exceed one hour. Our method can be integrated into the system without compromising the high level of asynchrony that is inherent in PoW blockchains.
We make several assumptions regarding attackers in our model. First, attackers are assumed to be capable of instantly becoming aware of any message that is publicly released on the network. In addition, they are assumed to have the ability to deliver any message to any miner instantaneously.
Figure 1: Overview of shared parital PoW.
@@ IV-B Details
[…]
The proposed method requires each miner to manage partial PoWs independent from the blockchain. It necessitates embedding new information, termed shared partial PoW, into blocks. We elaborate on how to embed shared partial PoW into a block in the case of Bitcoin in Section VI-A. Shared partial PoW is defined as the exclusive logical XOR of the hash values of the partial PoWs that belong to the set of valid partial PoWs for which $2\Delta_{B}+\Delta_{P}$ has elapsed since their arrival (see Figure 1). In addition, when propagating a block, miners must send the block as well as the set corresponding to the shared partial PoW. Hereafter, we refer to the exclusive logical sum of partial PoWs in the set, the set corresponding to the shared partial PoW, and partial PoWs that are valid and for which $2\Delta_{B}+\Delta_{P}$ has elapsed since their arrival as shared partial PoW.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV-C Parameter Settings
Considering the above theorem, the acceptance window $w$ is set to $\Delta_{B}$ in the proposed method.
@@
Considering the above theorem and the acceptance window $w=\Delta_{B}$, a valid partial PoW is considered as sufficiently shared $2\Delta_{B}$ after its arrival in the proposed method.
@@
Considering the above theorems and the acceptance window $w=\Delta_{P}$, our proposed method includes a valid partial PoW in the shared partial PoW $2\Delta_{B}+\Delta_{P}$ after its arrival.
@@ V Evaluation
In this section, we evaluate the effectiveness of our proposed method. In Section V-A, we demonstrate how effectiveness of our method in suppressing Match against post-generated block. This section includes both theoretical and simulation analyses.
In Section V-B, we evaluate the proposed method using ESM, which is a straightforward extension of SM that also considers Match against pre-generated block. We investigated the difference in the relative revenue for attackers between ESM and SM by varying the acceptance window. The analysis results demonstrate the importance of setting the acceptance window for a last-generated rule.

### v2

@@ Abstract
In the area of blockchain, numerous methods have been proposed for suppressing intentional forks by attackers more effectively than the random rule. However, all of them, except for the random rule, require major updates, rely on a trusted third party, or assume strong synchrony. Hence, it is challenging to apply these methods to existing systems such as Bitcoin.
To address these issues, we propose another countermeasure that can be easily applied to existing proof of work blockchain systems. Our method is a tie-breaking rule that uses partial proof of work, which does not function as a block, as a time standard with finer granularity. By using the characteristic of partial proof of work, the proposed method enables miners to choose the last-generated block in a chain tie, which suppresses intentional forks by attackers. Only weak synchrony, which is already met by existing systems such as Bitcoin, is required for effective functioning.
@@ I Introduction
In blockchain systems, transactions are processed in units known as blocks. Each block consists of a header and a list of transactions that are processed in that block. A block is valid when its header’s hash value is less than a certain target value. The process of identifying such a header is known as mining. The nodes that participate in the system and generate blocks are referred to as miners. When miners successfully generate a block, they publish it and the published block is shared across the network. Miners can receive rewards for the blocks that they successfully generate.
@@
The security of a blockchain system depends on each miner adhering to the aforementioned protocol. Therefore, it is desirable that following the protocol is the most profitable action for each miner. However, Eyal et al. introduced a new mining strategy known as selfish mining (SM) and demonstrated that blockchain does not always satisfy the necessary incentive compatibility[3]. Specifically, they demonstrated that miners can gain more rewards by withholding a successfully generated block instead of publishing it immediately, and then releasing it when another node publishes a block.
Eyal et al. proposed a tie-breaking rule that randomly selects a chain during chain ties (hereafter referred to as the random rule), thereby introducing the approach of modifying tie-breaking rules as a countermeasure to SM [3].
To suppress intentional forks more effectively than the random rule, several studies have been conducted[3, 4, 5, 6, 7, 8, 9, 10]. However, their solutions face the following three significant challenges: they necessitate major updates, require a trusted third party, or assume strong synchrony among miners.
To the best of our knowledge, Heilman was the first to propose a tie-breaking rule that selects the latest chain (hereafter referred to as last-generated rules)[4]. Last-generated rules enable each miner to distinguish between chains that are withheld by attackers and others. Consequently, last-generated rules can suppress intentional forks by attackers more effectively than the random rule.
@@
The proposed method is practical and applicable to existing systems that employ PoW blockchains, including Bitcoin. Our proposed method requires only that there exists a widely known upper bound on the message propagation times, and there exists a widely known upper bound on the absolute values of clock drifts. This synchrony is the minimum requirement for last-generated rules. Our approach does not require a specific trusted third party. Furthermore, it does not require updates that can destabilize the system, such as soft forks or hard forks.
@@
Organizations. The remainder of this paper is organized as follows. Section II describes intentional forks by attackers. Section III discusses related works. Section IV describes the proposed method. Section V demonstrates the effectiveness of the proposed method. Section VI presents a detailed discussion, primarily focusing on the application of the proposed method and consideration of clock drift. Section VII concludes the paper.
@@ II Intentional Forks by Attackers
In a blockchain, a fork refers to a situation in which a chain splits into multiple branches. There are two types of forks: those caused intentionally by miners not following the protocol and those that occur naturally. Attacks such as SM[3, 8, 11, 12, 13] are typical examples of intentional fork exploitation. Among intentional forks, chain ties that are intentionally created by attackers are known as intentional chain ties. SM and similar attacks exploit these intentional chain ties. The proposed method is a tie-breaking rule that aims to suppress intentional chain ties. Suppressing intentional chain ties leads to suppressing intentional forks, which in turn helps to mitigate attacks such as SM.
We consider SM as an example to demonstrate the exploitation of intentional forks. First, the proportion of a miner’s own blocks in the main chain is defined as the relative revenue. The goal of SM is to achieve a higher relative revenue than the proportion of the attacker’s hashrate in the total network. If this is accomplished, the attack is considered successful. The attacker withholds the block that they have successfully mined instead of publishing it immediately. When an honest miner generates a block, the attacker publishes their block, creating a chain tie. If a block is generated that extends the attacker’s chain after the chain tie, the honest miner’s block is invalidated, and vice versa. If the attacker mines blocks consecutively, they publish their block when the length difference between their chain and the honest miner’s chain is one. In this scenario, the honest miner’s chain is invalidated because the attacker’s chain is longer. Thus, the attacker can invalidate the blocks that are created by honest miners and increase their relative revenue by manipulating the timing of publishing their mined blocks in a manner that does not follow the protocol.
@@
Match against pre-generated block is difficult under the first-seen rule, however, it becomes feasible under last-generated rules. Last-generated rules impose a time limit on accepting chains in a tie to address this issue. This time limit is referred to as the acceptance window $w$. Without such a limit, attackers can always initiate chain ties against honest miners and invalidate their blocks. Conversely, Match against pre-generated block can be mitigated to an extent by setting this limit. We detail the extent to which setting an acceptance window quantitatively suppresses Match against pre-generated block in Section V-B.
@@ III Related Work
All of the aforementioned methods are tie-breaking rules. Zhang et al. proposed a fork choice rule known as weighted FRP to suppress SM[6]. This method involves including uncle blocks in a block to consider the number of published blocks in the main chain selection. However, this approach requires changes to the fork choice rule, which may be expensive to implement and can result in multiple forks. In addition, this method may lead to prolonged forks in situations in which the network is fragmented.
Zhang et al. also proposed another fork choice rule against SM[15]. The method uses in-blocks similar to partial PoWs. However, the method needs a soft fork and suppresses only a specific behavior of an attacker. That limits the effectiveness of the method.
Solat et al. proposed a method named ZeroBlock in which empty blocks are generated when no block is produced for a certain period, thereby invalidating withheld blocks[7]. However, this method requires precise clock synchronization across the entire system, which is a stringent requirement, especially for global-scale decentralized systems such as Bitcoin. Moreover, this approach assumes that blocks in PoW blockchains are generated at regular intervals, but the interval between block generations follows an exponential distribution in actual systems, which would limit the effectiveness of their proposed method. Additionally, implementing this method requires significant updates, such as hard forks, which makes it challenging to apply to systems that are already operational (e.g., Bitcoin).
Saad et al. proposed a fork choice rule with a similar concept to that of Lee et al.[8]. Their method uses high accuracy in predicting the block height, which is the time at which transactions are likely to be included in a block. Similar to the method of Lee et al., the average predicted processing times of transactions that are included in a block are used to distinguish between blocks generated by attackers and honest miners. However, their method also does not address the possibility of attackers including arbitrary transactions in their blocks. Furthermore, it applies only to specific patterns of attacker behavior, which limits its effectiveness.
@@
Sun et al. proposed a method that allows honest miners to delay block propagation to a selfish miner[10]. The method assumes that the network adopts the first-seen rule. The method reduces the winning rate of the attacker by delaying notification of the new block generated by honest miners. However, the method overlooks the following two critical issues. Firstly, they assume that all honest miners adopt the method. Even if one of the honest miners does not adopt the method, the effectiveness of the method is reduced. This is because honest miners who do not adopt the method may send blocks to the attacker as usual. Secondly, they assume that the blockchain network consists solely of mining pools. In reality, numerous solo miners are also participating in the network. The attacker can exploit this by setting up nodes to facilitate block propagation to itself.
@@
Figure 1: Blockchain functions as a clock that progresses block by block. Partial PoWs are easier to generate than blocks. Hence, the number of partial PoWs is larger than that of blocks. Consequently, the blockchain functions as a more precise clock by using partial PoWs.
@@ IV Proposed Method
Our proposed method is a type of last-generated rule. We use partial PoW, which originally had no functional use as a block, as a new time standard. Partial PoWs are easier to generate, can be produced in larger quantities than in blocks, and thus can function as a more granular clock (Figure 1).
The proposed method is practical. Widespread sharing of partial PoWs across the system is required for our method to be effective. However, a partial PoW typically has a smaller size than blocks and does not consume significant network resources from miners. In addition, the implementation of our method has a negligible impact on the transaction-processing capacity of the system. Our method does not require updates that could divide the hashrate of the system, such as hard forks or soft forks, nor does it require a trusted third party. Although a certain level of synchrony in the message propagation times and the progression of miners’ clocks (clock drift) are necessary, these requirements are achievable in most blockchain systems and do not compromise the high level of asynchrony that is inherent in PoW blockchains. These factors make our method more practical than previous methods.
Due to its practicality, the proposed method can easily be applied to existing systems such as Bitcoin. In addition, the proposed method is a tie-breaking rule, which provides the following two advantages. Firstly, the proposed method does not generate stale blocks resulting from its partial application. Secondly, the proposed method can be applied to existing fork choice rules. For example, it can be applied to rules such as GHOST[17], which had been partially adopted in Ethereum[14], and weighted FRP[6].
@@ IV-A Our Model
The final assumption is that the impact of clock drift among miners is sufficiently small to be negligible. For simplicity, We adopt this assumption in our explanation. However, this assumption can easily be replaced with a weaker one: there exists a widely known value $D$ such that the absolute value of the clock drift of any miner is less than $D$. We elaborate on how clock drift affects the proposed method in Section VI-D. In Bitcoin, the PoW blockchain ensures functionality provided that the clock discrepancies of miners do not exceed one hour. Our method can be integrated into the system without compromising the high level of asynchrony that is inherent in PoW blockchains. This is because we make assumptions only about clock drift, not about clock offset.
The above assumptions that there exists a widely known upper bound on the message propagation times, and there exists a widely known upper bound on the absolute values of clock drifts are the minimum requirement for any last-generated rule. This is because last-generated rules have an acceptance window and need to measure how much time has passed since the blocks arrived.
We make several assumptions regarding attackers in our model. First, attackers are assumed to be capable of instantly becoming aware of any message that is publicly released on the network. In addition, they are assumed to have the ability to deliver any message to any miner instantaneously. Our assumption allows attackers to detect and propagate blocks in any way.
Figure 2: How to include shared partial PoW in a block. Each miner considers a partial PoW “sufficiently shared” if $2\Delta_{B}+\Delta_{P}$ has elapsed since its arrival. Each miner includes sufficiently shared partial PoWs into a block and mines.
@@ IV-B Details
Here is the method with the fixed difficulty adjuster (hereafter referred to as the n-fixed method). We elaborate on the method with the variable difficulty adjuster (hereafter referred to as the n-variable method) in Section VI-A.
Partial PoW is a block header with a generation difficulty that is $1/n$ times lower than that of a valid block, where $n$ is the difficulty adjuster. That is, a block header $h$ functions as a valid block header when it satisfies $H(h)<\textit{Target}$, where $H$ is a cryptographic hash function and Target is a value that is determined by the difficulty. However, it also functions as a partial PoW if it satisfies $H(h)<n\textit{Target}$. Upon successfully generating partial PoW, miners publish and share it on the network, regardless of its validity as a block. As opposed to the propagation of transactions, the risk of DoS attacks owing to the propagation of partial PoWs is not a concern. This is because the validity of a partial PoW is verified, thereby providing proof of work[19].
The proposed method requires each miner to manage partial PoWs independent of the blockchain. It necessitates embedding new information, termed shared partial PoW, into blocks. We elaborate on how to embed shared partial PoW into a block in the case of Bitcoin in Section VI-A. Shared partial PoW is defined as the exclusive logical XOR of the hash values of the partial PoWs that belong to the set of valid partial PoWs for which $2\Delta_{B}+\Delta_{P}$ has elapsed since their arrival (see Figure 2). In addition, when propagating a block, miners must send the block as well as the set corresponding to the shared partial PoW. Hereafter, we refer to the exclusive logical sum of partial PoWs in the set, the set corresponding to the shared partial PoW, and partial PoWs that are valid and for which $2\Delta_{B}+\Delta_{P}$ has elapsed since their arrival as shared partial PoW.
@@
Algorithm 1 is an algorithm that applies our proposed method to the fork choice rule. The getMainChain function (line 1) uses multiple chains as arguments and determines one main chain. In line 2, the fork choice rule (forkChoiceRule function) is used to select the main chain candidates. Subsequently, the earliest arrival time ($earliestTime$) among the main chain candidates is determined up to line 6. Next, from the main chain candidates, those whose arrival time is within $w$ from $earliestTime$ are selected (line 9), and among these, a main chain is uniquely determined using the proposed method (line 10). The chainMax function (line 16) uses two chains as arguments and returns the one that contains more shared partial PoWs. Line 17 determines whether $\textit{C}_{\textit{1}}$ is an empty chain. Thereafter, the chainWeight function (line 31) determines which chain contains more shared partial PoWs and returns the chain with more shared partial PoWs. If the number is equal, one chain is randomly selected. The chainWeight function (line 31) uses a single chain as an argument and returns the number of shared partial PoWs in that chain. First, each block belonging to the chain is selected to obtain the total shared partial PoW of the entire chain (lines 32-35). Subsequently, each shared partial PoW belonging to the chain is selected. If it is valid and sufficiently shared (line 38), 1 is added to res (line 39); otherwise, $-1$ is returned (line 41). The validity of the shared partial PoW was verified using the valid function. Finally, res is returned (line 44).
@@
1:procedure getMainChain($\textit{C}_{\textit{1}},\ldots,\textit{C}_{\textit{k}}$)
2:     $\textit{candidates}\leftarrow\text{forkChoiceRule}(\textit{C}_{\textit{1}},\ldots,\textit{C}_{\textit{k}})$
3:     $\textit{earliestTime}\leftarrow 0$
@@
16:procedure chainMax($\textit{C}_{\textit{1}},\textit{C}_{\textit{2}}$)
17:     if $\textit{C}_{\textit{2}}=\epsilon$ then
18:         return $\textit{C}_{\textit{1}}$
@@
20:     if $\text{chainWeight}(\textit{C}_{\textit{1}})>\text{chainWeight}(\textit{C}_{\textit{2}})$ then
21:         return $\textit{C}_{\textit{1}}$
@@
23:         if $\text{chainWeight}(\textit{C}_{\textit{1}})<\text{chainWeight}(\textit{C}_{\textit{2}})$ then
24:              return $\textit{C}_{\textit{2}}$
@@
26:              return $\textit{C}_{\textit{1}}$ or $\textit{C}_{\textit{2}}$ randomly
@@ IV-C Parameter Settings
Considering the aforementioned theorem, the acceptance window $w$ is set to $\Delta_{B}$ in the proposed method.
@@
Considering the aforementioned theorem and the acceptance window $w=\Delta_{B}$, a valid partial PoW is considered as sufficiently shared $2\Delta_{B}$ after it arrives in the proposed method.
@@
Considering the aforementioned theorems and the acceptance window $w=\Delta_{P}$, our proposed method includes a valid partial PoW in the shared partial PoW $2\Delta_{B}+\Delta_{P}$ after its arrival.
@@ V Evaluation
In this section, we evaluate the effectiveness of our proposed method. In Section V-A, we demonstrate the effectiveness of our method in suppressing Match against post-generated block. This section includes both theoretical and simulation analyses.
In Section V-B, we evaluate the proposed method by using ESM, which is a straightforward extension of SM that also considers Match against pre-generated block. We investigated the difference in the relative revenue for attackers between ESM and SM by varying the acceptance window. The analysis results demonstrate the importance of setting the acceptance window for a last-generated rule.
(a) Situation that honest miners even.
(b) Situation that honest miners win.
(c) Situation that honest miners may lose.
Figure 3: If the attacker successfully generates a block, they must publish their partial PoW. This is because if an insufficiently shared PoW is included in a block, honest miners will assign a value of $-1$ to the block, leading to the attacker’s block not being chosen in the event of a tie. It takes $2\Delta_{B}+2\Delta_{P}$ for honest miners to include one partial PoW after it has been published. This duration includes $\Delta_{P}$ for the partial PoW to be propagated across the network and $2\Delta_{B}+\Delta_{P}$ for each miner to consider it sufficiently shared and include it in a block. Figure 3(a) shows the situation in which an honest miner’s block is chosen randomly in a tie. This is because the attacker’s shared partial PoWs that have been published are included in the block generated by the honest miner. Figure 3(b) shows the situation in which an honest miner’s block is chosen in a tie. Figure 3(c) shows the situation in which an honest miner’s block may not be chosen in a tie. This is because the attacker’s partial PoW that has been published may not be included in blocks honest miners are mining.